# Li He1, Pei-Yang Lin1, Peng-Hu Wei5, Guo-Guang Zhao5, Yun-Zhe Liu†,1,4 1

### v1

@@ 15. Answer: [Yes]
  16. Justification: We provide detailed information related to our model and baselines in Appendix C and Appendix B, respectively. Besides, we provide code and demo dataset in Section 6.
@@ 19. Answer: [Yes]
  20. Justification: We provide code and demo dataset in Section 6. Due to the lack of open-source sEEG datasets related to language, we collected a well-annotated Chinese word-reading sEEG dataset, and evaluated our model on this dataset. However, respecting the efforts of the data collectors, we only provide the dataset of some subjects to reproduce the experimental results in the main text. The whole dataset will be publicly available to ensure the reproducibility of this work.

### v3

@@ 15. Answer: [Yes]
  16. Justification: We provide detailed information related to our model and baselines in Appendix C and Appendix B, respectively. Besides, we provide code and dataset in https://github.com/liulab-repository/Du-IN.
@@ 19. Answer: [Yes]
  20. Justification: We provide code and dataset in https://github.com/liulab-repository/Du-IN. Due to the lack of open-source sEEG datasets related to language, we collected a well-annotated Chinese word-reading sEEG dataset, and evaluated our model on this dataset. However, respecting the efforts of the data collectors, we only provide the dataset of some subjects to reproduce the experimental results in the main text. The whole dataset will be publicly available to ensure the reproducibility of this work.